OpenStudy (anonymous):

Choose the true statement about the graph of x < 4. closed circle on 4 and all numbers to the left shaded open circle on 4 and all numbers to the right shaded open circle on 4 and all numbers to the left shaded

OpenStudy (anonymous):

closed circle on 4 and all numbers to the left shaded

OpenStudy (usukidoll):

while the left shaded part is true. If we have a > or < sign, it should be an open circle

OpenStudy (usukidoll):

should we have a \[\leq \] or \[\geq \] we have a closed circle. I don't see those signs for x<4

OpenStudy (anonymous):

and if they give that like this Choose the true statement about the graph of 3x ≥ 6. closed circle on 2 and all numbers to the right shaded closed circle on 2 and all numbers to the left shaded open circle on 2 and all numbers to the right shaded open circle on 2 and all numbers to the left shaded

OpenStudy (usukidoll):

for that question, solve the inequality first. then since you have a \[\geq \] we have a closed circle

OpenStudy (usukidoll):

and graph according to the point of the symbol

OpenStudy (usukidoll):

for example if I want to graph x < 2 on a number line 

OpenStudy (anonymous):

so it will be the second one

OpenStudy (usukidoll):